OpenStudy (anonymous):

A ball of mass m moving at speed v0 collides head on with a stationary ball of mass 3m. If the collision is completely inelastic, what is the ratio Kf Ki of the kinetic energies of the balls before and after the collision? 1. 1/2 2. 1/9 3. 1/3 4. 1, since Ki = Kf for any collision in which no external forces act. 5. 4/16 6. 1/4 7. Zero, since a completely inelastic collision does not conserve energy. 8. 9/16 9. 1/16 10. 2/9

OpenStudy (loser66):

@dan815 I don't know why, my answer is not one of them. before collision, K_i = 1/2 mv^2 after collision, K_f = 1/2 (m +3m) v^2 = 2mv^2 so, the ratio K_f/K-i =4 which is not one of the choices. why? what's wrong with my logic?

OpenStudy (anonymous):

not sure... but the answer is 1/4 :)